# Sparse Autoencoders Learn Monosemantic Features in Vision-Language Models

### v2

@@ 4.3 Steering Multimodal LLMs
Figure 6 illustrates the effectiveness of concept insertion by manipulating a single neuron of the SAE. We prompt the model with the instruction “Write me a short love poem”, along with a white image. By intervening on an SAE neuron associated to the pencil concept and increasing the corresponding activation value, we observe the impact on the generated output text. While the initial output mentions the “white” color and focuses on the textual instruction, i.e. “love” poem, the output becomes more and more focused on pencil attributes as we manually increase the intervention value $\alpha$ (most highly activating images for the selected neuron is in Appendix) until it only mentions pencils. We provide more examples with a different input prompt (“Generate a scientific article title”) in Appendix, for which steered LLaVA exhibits a similar behavior.
To further confirm the steering capabilities, we prompt the model to answer “What is shown on the image? Use exactly one word.” and compare its original output with the one generated after fixing a specific SAE neuron activation value at $\alpha=100$, one neuron after the other. More specifically, we measure embedding similarity in CLIP space between the output word and the top-$16$ images activating the selected neuron. In the first setup, we input the same white image as above, and intervene on the first $1000$ neurons. In the second setup, we use $1000$ random images from ImageNet as inputs and steer only the $10$ first neurons.
@@
Table 2: Average similarity between neurons’ highly activating images and output word, with and without steering, using a white input image or random images from ImageNet. Upper-bound score with correct image and class name text pairs is $0.283\pm 0.034$, while lower-bound with random image and class name text pair is $0.185\pm 0.028$.
@@
We report results for the two setups in Table 2. In both setups, steering SAE neurons induces higher similarity between images highly activating the imputed neuron and the modified text outputs, compared to the original text outputs. To put these results in perspective, we additionally measure the average similarity scores between one representative image of each ImageNet class with the class name text as an upper-bound value, and the average scores between randomly selected pairs of image and class name text as a lower-bound value. For the lower-bound, we obtain an average score of $0.185\pm 0.028$ and $0.283\pm 0.034$ for the upper-bound. These results illustrate the significance of the relative improvement we obtain from steering, i.e. a relative improvement of about $22\%$ when considering the range between the possible lower- and upper-bounds.
@@ 5 Conclusion
Limitations. We primarily focused on SAEs as tools to explain VLMs. Investigating and comparing monosemanticity of other forms of dictionary learning [32] could further strengthen our findings.
@@ Appendix C User study
Alignment $\text{Alignment Score}=\frac{1}{|\mathcal{Q}|}\sum_{(k_{x},k_{y})\in\mathcal{Q}}\delta_{(k_{x},k_{y})},$
@@
In total, we collected 1,000 user pair rankings with the help of 71 annotators on the Mechanical Turk platform. The number of answers per annotator ranged from 1 to 205, with a median of 24. Annotators were compensated at a rate of \$$\$0.02$ per answer.
@@ E.1 Unnormalized plots
Figure A6: MS in decreasing order across neurons. Results are shown for a layer without SAE (“No SAE”), and with SAE using different expansion factors (×1,×2,×4,×16$\times$ 1$\times$ 2$\times$ 4$\times$ 16× 1 , × 2 , × 4 , × 16 and ×$\times 64$).
@@ E.2 Detailed statistics and more models
Figure A7: MS in decreasing order across neurons. Results are shown for the last layers of two different models, without SAE (black dashed line), and with SAE being trained with expansion factor $1$ ( green solid line). MS is computed with distinct image encoders $E$.
@@
Figure A9: Monosemanticity Scores (computed using CLIP ViT-B) in decreasing order across neurons without normalizing by width. Results are shown for a layer without SAE (“No SAE”), and with SAE using different expansion factors (×1,×4$\times$ 1$\times$ 4× 1 , × 4 and ×$\times 16$).
@@ E.3 Matryoshka hierarchies
We train and evaluate the SAE on embeddings extracted from iNaturalist [41] dataset using an expansion factor $\varepsilon=2$ and groups of size =$\mathcal{M}=\{3,16,69,359,1536\}$. These group sizes correspond to the numbers of nodes of the first 5 levels of the species taxonomy tree of the dataset, i.e. the respective number of “kingdoms”, “phylums”, “classes”, “orders”, and “families”.
@@ Appendix F Reconstruction of SAEs
[…]
@@ Appendix G Uniqueness of concepts
The sparse reconstruction objective regularizes the SAE activations to focus on different concepts. To confirm it in practice, we collect top-16 highest activating images for each neuron of SAE and compute Jaccard Index $J$ between every pair of neurons. The images come from training set. We exclude $10$ out of $12288$ neurons for which we found less than $16$ activating images and use Matryoshka SAE trained on the last layer with expansion factor of $16$. We find that $J>0$ for $16000$ out of $75368503$ pairs ($>0.03\%$) and $J>0.5$ for only $20$ pairs, which shows very high uniqueness of learned concepts.

### v3

@@ 4.3 Steering Multimodal LLMs
Quantitative Results. We first compare the performance of SAE-based steering for VLMs against Difference-in-Means (DiffMean) [2], a popular approach based on activation steering. For each of $100$ SAE neurons in LLaVA, we identify its top-activating image grid. To perform concept insertion, we boost the neuron’s activation and prompt LLaVA with $10$ diverse text queries, such as “Propose a math word problem,” or “Invent a new holiday,” then evaluate if the output contains the concept and still responds to the prompt. For concept supression, we apply a negative intervention, ask LLaVA to describe the images, and check whether the concept is removed and unrelated images are still described correctly. The evaluation is done with a LLM-as-a-judge setup using GPT-4.1-mini [43]. We provide the full details of the prompts and the evaluation procedure in Appendix B. The results in Table 2 demonstrate that SAE-based directions outperform DiffMean in both concept insertion and suppression. For insertion, SAE effectively introduces the intended concept ($48.7\%$ vs. $53.1\%$) while maintaining much stronger adherence to the base prompt ($85.5\%$ vs. $66.2\%$). For suppression, it performs on par with DiffMean in removing the target concept ($64.4\%$ vs. $64.0\%$), yet far surpasses it in preserving unrelated content ($81.4\%$ vs. $38.7\%$).
@@
Table 3: Mean similarity of neurons’ activating images to output word, with and without steering, on white or random ImageNet images. Upper bound with correct image-classname pairs is $0.283$, lower bound with random pairs is $0.185$.
@@
To further confirm the steering capabilities of SAE neurons, we also perform a sanity check on the steering capabilities of the SAE neurons. We directly compare CLIP similarity scores between the top-$16$ images activating a given neuron and the corresponding text outputs, both before and after steering. We prompt the model with “What is shown on the image? Use exactly one word.” and compare its original answer to the one generated after fixing the activation of a specific SAE neuron to $\alpha=100$, varying one neuron at a time. In the first setup, a white image is used while intervening on the first 1000 neurons to isolate the neuron manipulation’s effect. In the second, 1000 random ImageNet images are used while steering only 10 neurons to test effects on natural inputs. The results in Table 3 clearly illustrates that steering increases image-text similarity scores. For context, we compute reference similarities in CLIP space: the average similarity between each ImageNet class image and its class name sets an upper bound ($0.283\pm 0.034$), while random image–class pairs set a lower bound ($0.185\pm 0.028$). Neuron steering yields a relative gain of $22\%$ within this range, highlighting the significance of the results.
Qualitative Examples. Figure 6 illustrates the effectiveness of concept insertion by manipulating a single neuron of the SAE. We prompt the model with the instruction “Write me a short love poem”, along with a white image. By intervening on an SAE neuron associated to the pencil concept and increasing the corresponding activation value, we observe the impact on the generated output text. While the initial output mentions the “white” color and focuses on the textual instruction, i.e. “love” poem, the output becomes more and more focused on pencil attributes as we manually increase the intervention value $\alpha$ (most highly activating images for the selected neuron is in Appendix) until it only mentions pencils. We provide more examples with a different input prompt (“Generate a scientific article title”) in Appendix, for which steered LLaVA exhibits a similar behavior.
@@ 5 Conclusion
Limitations. We focused our evaluation on various SAE architectures, the most common dictionary learning implementations that scale effectively to large VLMs. However, our MS metric is model-agnostic and could be applied to study other comparable approaches as well. High MS score neurons do not always produce precise effects when used to steer MLLM outputs. For example, a golden retriever neuron from an SAE trained on ImageNet can trigger any dog-related output. This could happen because, while SAEs can disentangle detailed classes in the dataset, MLLMs may have limited fine-grained understanding and may not be perfectly aligned with the vision encoder. Moreover, a fraction of the SAE neurons that act as feature detectors do not exhibit any clear steering effect [1].
@@ Appendix B More details on steering
The steering capabilities discussed in Section 4.3 are evaluated using an LLM-as-judge setup with the following prompts:
  • “Write me a short love poem,”
  • “Generate a scientific article title,”
  • “Give me a four-item to-do list,”
  • “Write me a two-verse rap song,”
  • “Propose a math word problem,”
  • “Write a paragraph from a Wikipedia page,”
  • “Invent a new holiday,”
  • “Write a dialogue,”
  • “Write a newspaper headline and first paragraph,”
  • “Give a conversation starter for a party.”
@@ Appendix C User study
$\text{Alignment Score}=\frac{1}{|\mathcal{Q}|}\sum_{(k_{x},k_{y})\in\mathcal{Q}}\delta_{(k_{x},k_{y})},$
@@
In total, we collected 1,000 user pair rankings with the help of 71 annotators on the Mechanical Turk platform. The number of answers per annotator ranged from 1 to 205, with a median of 24. Annotators were compensated at a rate of $\mathdollar 0.02$ per answer.
@@ E.1 Unnormalized plots
Figure A6: MS in decreasing order across neurons. Results are shown for a layer without SAE (“No SAE”), and with SAE using different expansion factors ($\times 1,\times 2,\times 4,\times 16$ and $\times 64$).
@@ E.2 Detailed statistics and more models
Figure A7: MS in decreasing order across neurons. Results are shown for the last layers of two different models, without SAE (black dashed line), and with SAE being trained with expansion factor 11 ( green solid line). MS is computed with distinct image encoders $E$.
@@
Figure A9: Monosemanticity Scores (computed using CLIP ViT-B) in decreasing order across neurons without normalizing by width. Results are shown for a layer without SAE (“No SAE”), and with SAE using different expansion factors ($\times 1,\times 4$ and $\times 16$).
@@ E.3 Matryoshka hierarchies
We train and evaluate the SAE on embeddings extracted from iNaturalist [55] dataset using an expansion factor $\varepsilon=2$ and groups of size $\mathcal{M}=\{3,16,69,359,1536\}$. These group sizes correspond to the numbers of nodes of the first 5 levels of the species taxonomy tree of the dataset, i.e. the respective number of “kingdoms”, “phylums”, “classes”, “orders”, and “families”.
@@ Appendix F Reconstruction of SAEs
Table A10: Comparison of true sparsity measured by $L_{0}\text{-norm}$ for different SAEs trained with $K=20$ for a CLIP ViT-L model.
@@ Appendix G Uniqueness of concepts
The sparse reconstruction objective regularizes the SAE activations to focus on different concepts. To confirm it in practice, we collect top-16 highest activating images for each neuron of SAE and compute Jaccard Index $J$ between every pair of neurons. The images come from training set. We exclude $10$ out of $12288$ neurons for which we found less than $16$ activating images and use Matryoshka SAE trained on the last layer with expansion factor of $16$. We find that $J>0$ for $16000$ out of $75368503$ pairs ($0.03\%$) and $J>0.5$ for only $20$ pairs, which shows very high uniqueness of learned concepts.